A ball is thrown upward. It leaves the hand with a velocity of 13.8 m/s, having been accelerated through a distance of 0.460 m. Compute the ball's upward acceleration, assuming it to be constant.
Answer #1

here,

the final velocity , v = 13.8 m/s

s = 0.46 m

let the accelration be a

initial velocity , v0 = 0 m/s

using third equation of motion

v^2 - v0^2 = 2 * a * s

13.8^2 - 0 = 2 * a * 0.46

solving for a

a = 207 m/s^2

the accelration is 207 m/s^2